A circular clock has the numbers 1 through 24 evenly spaced around it. Which of the following answers include the numbers that is directly across from 7?
A 16 or 17 B 18 or 19 C 20 or 21 D 5 E 2

Answer by Edwin McCravy(20055)

A circular clock has the numbers 1 through 24 evenly spaced around it. Which of the following answers include the numbers that is directly across from 7?
A 16 or 17 B 18 or 19 C 20 or 21 D 5 E 2

I can't understand the word "numbers" (plural) instead of just "number"
(singular).  There is only ONE number opposite each number, not TWO numbers!

On an ordinary 12-hour clock, 12 is opposite 6, 1 is opposite 7, 2 is
opposite 8, etc., n is opposite n+6, for n= 1,..,6, and n is opposite
n-6 for n=7,...12.

Similarly on a 24-hour clock, 24 is opposite 12, 1 is opposite 13, 2 is
opposite 14, etc., n is opposite n+12, for n = 1,..,12, and n is opposite
n-12 for n=13,..,24.

Therefore on a 24-hour clock, 7 is opposite 7+12 or 19.

So B is the only choice of two numbers of which one of them is opposite the
number 7.  But you should ask your teacher about why the problem seems to
imply that there are two numbers directly across from each number, when there
is really only one number directly across from each number.     

Edwin